OpenStudy (anonymous):

Hi everyone, im having trouble with this question, if you could help me out i would very grateful. The question says the point (3,-4) is on the graph y=f(x). What would be the new point on the transformed graph y-5=2f(3(x-1))???

OpenStudy (anonymous):

There's not enough info... we need to know more about f(x). For more solutions to textbook problems, be sure to check out http://www.slader.com/s/eWFrZXlnbGVl. I submit a lot of content there as well. (They even pay you if you help them out to add solutions).

OpenStudy (anonymous):

you can solve \[3(x-1)=3\] for x and get \[3x-3=3\] \[3x=6\] \[x=2\]

OpenStudy (anonymous):

so if \[x=2\] then \[3(x-1)=3\] so if \[x=2\] you get \[f(3(x-1))=f(3)=-4\]and then \[2\times -4+5=-3\] so the new point on the graph will be \[(2,-3)\]

OpenStudy (anonymous):

Ok that makes way more sense! Thanks so much!!

OpenStudy (anonymous):

yw